### v1

@@ 4.3 Control System Overview and Implementation
The overall control of the developed system includes five main components (Fig. 2). Manipulation strategy transitions are managed by predefined task objectives. This system employs a state machine that dynamically alters between three modes: translation, rotation, and flipping. The position and the orientation of the object are obtained by the real-time sensor data from the vision system, guiding the dynamic generation of the surfaces’ roll, pitch, and height references in accordance with the appropriate manipulation strategy. Using the inverse kinematics algorithm explained in Method S1, the desired joint angles are computed. These reference values are then input into a PID-based servo motor controller, ensuring precise trajectory tracking of the joint commands. In the control scheme for the BLDC motor, as illustrated in Fig. 2c, a cascade control structure is implemented, comprising both a feedforward component and a current loop within the position loop.
@@
2:     ##\#\ #Line 3-4: compute the position error between the target and the object’s current position.
@@
6:         ##\#\ #Line 7-10: if the positional error exceeds $\epsilon$, calculate the reference pose for the surface.
7:         $\theta_{surface\_roll}=PID(Y_{error})$triangleright Equation 2
8:         $\theta_{surface\_pitch}=PID(X_{error})$triangleright Equation 2
9:         $H=Asin(\omega t)+H_{init}$triangleright Equation 3
@@
[…]
In low-friction scenarios, where the friction coefficient $\mu$ is smaller than $tan(\psi)$ (the tangent of the surface’s normal vector polar angle $\psi$), the control approach resembles a classic ball balancing table. Here, adjusting the surface’s inclination smoothly guides the object to its target. Under higher friction, inclination alone may not suffice. To overcome this, we modulate the surface’s height as well as its orientation (Fig. 2b), introducing vertical oscillations that reduce frictional resistance and apply periodic impulses to move the object. We define:
@@
wehere $t$ is the elapsed time and $\omega$ is the oscillation frequency, adjustable to suit different objects. The elevation of the surface also affects its maximum achievable tilt angle (see Fig. S8). To maintain a large range of motion, here the vertical displacement amplitude spans from 8 cm to 10 cm.
@@
$C_{WR}=\begin{bmatrix}&cos(\theta)\ &0\ &sin(\theta)\\ &sin(\phi)sin(\theta)\ &cos(\phi)\ &-cos(\theta)sin(\phi)\\ &-cos(\phi)sin(\theta)\ &sin(\phi)\ &cos(\phi)cos(\theta)\\ \end{bmatrix}.$ (4)
@@
$N_{O_{R}}=C_{WR}*[0,0,1]^{T}=[sin(\theta),-cos(\theta)sin(\phi),cos(\phi)cos(\theta)]^{T}.$ (5)
@@
$\begin{split}\delta&=arctan(\frac{N_{O_{R}}(y)}{N_{O_{R}}(x)})=arctan(\frac{-cos(\theta)sin(\phi)}{sin(\theta)}),\\ \psi&=arccos(N_{O_{R}}(z))=arccos(cos(\phi)cos(\theta)).\end{split}$ (6)
Once we determine the required orientation and height parameters $(\delta,\psi,H)$ for the Canfield mechanism’s end-effector, we apply inverse kinematics (Method S1, Chapter Method S1: Single Module System Modeling) to compute the corresponding joint angles.
@@
2:     ##\#\ #Line 3-4: initialize the surface.
@@
5:     ##\#\ #Line 6-21: repeat when the angular error between the object’s current angle and target angle exceeds $\epsilon$. This constitutes one cycle.
@@
7:         $\theta_{surface2\_roll}=PID(\theta_{ref}-\theta_{yaw})$triangleright Equation 7
@@
13:         ##\#\ #rotate the surface to the reference pose to induce the object’s rotation.
@@
[…]
17:         ##\#\ #reset the surface after one cycle to establish new contact points in next cycle.
@@
22:     ##\#\ #Line 23-24: reset the surface once the object reaches the reference angle to prepare for the next task.
@@
[…]
@@
2:     ##\#\ #Line 3-5: based on the object’s size, transport it from the initial position to a flipping position between the two surfaces.
@@
4:         $Surface_{1}\leftarrow transfer\left(X_{ref},Y_{ref}\right)$triangleright Algorithm 1
@@
7:         ##\#\ #Line 8-9: Lower the surface to initialize the flipping task.
@@
12:         ##\#\ #Quickly raise Surface 2 to generate an impact, flipping the object.
[…]
15:         ##\#\ #Line 16-17: after flipping, reset the surface to prepare for the next task.
@@
$\begin{split}0-\frac{1}{2}J\omega_{t1}^{2}&\geq\frac{L}{2}mg(sin(\theta)-1),\\ \omega_{t1\_min}&=\sqrt{\frac{mgL(1-sin(\theta))}{J}}.\end{split}$ (8)
@@ Acknowledgements
The authors thank Alexander Schüßler of Reconfigurable Robotics Lab, EPFL for his contributions to using computer vision to identify the position of deformable objects.
Funding
This research has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement no: 101069536 (MOZART project).
@@
Supplementary material
  • Method S1. Single Module System Modeling.
  • Method S2. Spherical Cordinate-based Workspace Evaluation.
  • Method S3. Multi-module Transfer Strategy.
  • Table S1. Notation for kinematics analysis.
  • Algorithm S1. Multi-module transition: Translate object from Surface 2 to 1.
  • Figure S1. Module design and workspace analysis.
  • Figure S2. Workspace analysis in spherical coordinates.
  • Movie S1. Overview of surface-based manipulation.
  • Movie S2. Validation of translation strategy with randomly shaped objects. A piece of randomly shaped putty is placed at different initial positions on the surface and translated to a designated target location in the upper-left corner. Six consecutive experiments were conducted, each starting from a different initial condition, to illustrate the performance of the translation strategy.
  • Movie S3. Validation of flipping strategy. The object initially placed at a random position on the surface, is translated to the space between two planes to perform a flipping task. This sequence is executed automatically and repeated two times, demonstrating the robustness and reliability of the flipping strategy.
  • Movie S4. Validation of rotation strategy. The object is rotated about its Z-axis to multiple target angles (20°, 30°, 60°, and 80°) in sequence. After achieving each target orientation, the system briefly maintains the object’s angle before proceeding to the next reference angle.
  • Movie S5. Shape manipulation of free-form objects. A long strip of putty undergoes two successive folds by combining multiple previously introduced manipulation strategies.
Method S1: Single Module System Modeling
We adopted a modular design that consists of identical units. Modeling a single module allows us to understand the workspace of each unit and optimize their control to accomplish manipulation tasks. Developing an analytical model for a single module also facilitates scalability in both size and number of modules, making it easier to design larger systems or integrate additional modules as needed. To understand the relationship between the end-effector position and the joint input angles, we studied the system’s kinematics. Our structure is based on an origami-inspired Canfield mechanism [52], whose kinematics have been explored in previous studies [54, 68, 69]. To accurately represent the kinematic equations of the structure illustrated in Figure S7(a), we provided a schematic representation in Figure S7(b). Notably, the water bomb component of the leg in Figure S7(a) functions as a spherical joint within a certain range of motion. Each physical parameter in the schematic is detailed in Table S2.
Table S2: Notation for kinematics analysis
Notion Description
$O_{R}\in\mathbb{R}^{3}$ Geometric center of the top surface
$R_{i}\in\mathbb{R}^{3}$ Revolute joint connecting the leg and the top surface
$M_{i}\in\mathbb{R}^{3}$ The midpoint of each origami leg
$O_{M}\in\mathbb{R}^{3}$ The center of the virtual plane formed by $M_{1}$, $M_{2}$, $M_{3}$
$B_{i}\in\mathbb{R}^{3}$ Revolute joint connecting the leg and the base surface
$O_{B}\in\mathbb{R}^{3}$ Geometric center of the base surface
$\theta_{i}\in\mathbb{R}^{3}$ The angle between the leg and the base plane
$l\in\mathbb{R}$ The length of $M_{i}B_{i}$, where $M_{i}B_{i}=M_{i}R_{i}$
$r\in\mathbb{R}$ The length of $O_{B}B_{i}$, where $O_{B}B_{1}=O_{B}B_{2}=O_{B}B_{3}$
$\phi_{i}\in\mathbb{R}$ The leg’s azimuth angle. $\phi_{1}=0,\phi_{2}=\frac{2\pi}{3},\phi_{3}=\frac{4\pi}{3}$
Forward Kinematics
To understand the relationship between the joint angles and the position and orientation of the end effector, and thereby determine its workspace, we studied the system’s forward kinematics.
Assuming $O_{B}=[0,0,0]^{T}$ as the origin of the world coordinate system, the geometric configuration allows us to compute the coordinates of the midpoint $M_{i}$ of each leg based on the input angles of the legs $\theta_{i}$:
$M_{i}=\begin{bmatrix}cos\phi_{i}(r+lcos\theta_{i})\\ sin\phi_{i}(r+lcos\theta_{i})\\ lsin\theta_{i}\end{bmatrix},$ (10)
where $\phi_{i}$ represents the leg’s azimuth angle, we can assume $\phi_{1}=0,\phi_{2}=\frac{2\pi}{3},\phi_{3}=\frac{4\pi}{3}$.
Then, utilizing the algorithm described in [54], we can obtain the analytical solution for the Cartesian coordinates of the center of the top surface $O_{R}$.
By substituting the actual parameters $l=5.9cm$, $r=4.4cm$, and constraining $\theta_{i}$ within the range of $[30\degree,90\degree]$, we can map out the workspace of the top plane in the Cartesian coordinate system as in Fig. S7(c). Here, the points’s positions represent all possible coordinates of the top plane’s center, with the highest center point reaching approximately $11.8cm$ and the lowest at about $5.9cm$. Due to the distinctiveness of the Canfield structure, this parallel mechanism is a 3 DoF pointing system. As such, when the X and Y coordinates of the center of the upper surface change, the tilt angle of the upper surface also varies [69]. Therefore, for our application, unlike the current studies [54, 69] pertaining to continuum robots, knowing the center position of the plane only is insufficient. To manipulate objects through the deformation of a surface composed of multiple modules, it is crucial for us to comprehend the relationship between the tilt angle of the end effector and the input leg angles. For this, we first calculate the normal vector of the virtual plane $M_{1}M_{2}M_{3}$:
$N=\left[N_{x},N_{y},N_{z}\right]=(M_{1}-M_{2})\times(M_{1}-M_{3}).$ (11)
Since the top plane $R$ and the base plane $B$ are symmetrical with respect to the virtual plane $M$ [69], we can determine the representation of the normal vector of the top plane in the spherical coordinate system. Here, $\psi$ denotes the polar angle between the normal vector of the top plane and the positive Z-axis, and $\delta$ signifies the azimuthal angle between the normal vector and the positive X-axis:
$\begin{split}\psi&=2arccos(\frac{N_{z}}{\sqrt{N_{x}^{2}+N_{y}^{2}+N_{z}^{2}}}),\\ \delta&=arctan(\frac{N_{y}}{N_{x}}).\\ \end{split}$ (12)
Inverse Kinematics
In the design shown in Fig. S7, the three motors that can perform closed-loop position control are abstracted as nodes $B_{i}$ in Fig. S7(b), with the $\theta$ representing the motor output angle. To facilitate the control of the robot for more complex closed-loop interactive behaviors in the task space, we need to understand the relationship between specific plane poses and the motor output angles. Therefore, we conducted a study on the inverse kinematics of the robot.
When describing the pose of the top square plane, we use the roll and pitch from XYZ-Euler angles (since the mechanism lacks the yaw DoF in XYZ-Euler angles) and the distance from the plane center to the base to characterize the plane’s distance, which could be noted as $[Roll,Pitch,H]$. In Equation (6) , we explore how to convert XYZ-Euler angles into the representation of the top plane’s normal vector in spherical coordinates. In this chapter, we will focus on discussing how to determine the motor angles corresponding to $[\delta_{top},\psi_{top},H]$. Since the top plane of the Canfield structure is symmetrical to the bottom plane about a central virtual plane, we can obtain the representation of the inclination angle of the normal vector of the virtual plane $M_{1}M_{2}M_{3}$ in spherical coordinates:
$\begin{split}\psi_{virtual}&=\frac{\psi_{top}}{2},\\ \delta_{virtual}&=\delta_{top}.\\ \end{split}$ (13)
The distance $r_{0}$ from the center of the base to the top is:
$r_{0}=\frac{H}{sin(\frac{\pi}{2}-\psi_{virtual})}.$ (14)
Finally, according to [54] desired legs angle $\theta_{i}$ could be derived from this second-order polynomial equation:
$\theta_{i}=2arctan(t_{i}),\theta_{i}\in[0,\frac{\pi}{2}],$ (15)
where:
$t_{i}=\frac{-b_{i}\pm\sqrt{b_{i}^{2}-4a_{i}c_{i}}}{2a_{i}},$ (16)
$\begin{split}&a_{i}=(r-l)(sin(\psi_{virtual})cos(\delta_{virtual}-\theta_{i}))-\frac{r_{0}}{2},\\ &b_{i}=2lcos(\psi_{virtual}),\\ &c_{i}=(r+l)(sin(\psi_{virtual})cos(\delta_{virtual}-\theta_{i}))-\frac{r_{0}}{2}.\\ \end{split}$ (17)
Dynamics
To optimize the manipulation capabilities of our system, it is crucial to understand how the robot’s end effector exerts forces at different positions within its workspace under static conditions. Analyzing the variation in the maximum external force output allows us to select the most appropriate operational workspace for manipulation tasks. We then model the robot’s equations of motion (EoM) using projected Newton-Euler (PNE):
$M(q)\ddot{q}+b(q,\dot{q})+g(q)=\tau+J_{P}^{T}F$ (18)
Where $q$ represents the three motor angles, $M(q)$ is the mass matrix, $b(q,\dot{q})$ is the coriolis and centrifugal terms, g(q) is gravitational terms, $\tau$ is the generalized torques of generalized coordinates, $F$ is the external force acting on the top surface, and $J_{P}$ is the translational Jacobian of the top plane’s geometric center.
The calculation of translational Jacobian J can be obtained by taking the partial derivative for the leg joint angle, based on the relationship between the planar center position $O_{R}\in\mathbb{R}^{3}$ and the motor angle $\theta\in\mathbb{R}^{3}$in the forward kinematics equation.
$J_{P}=\frac{\partial O_{R}(\theta)}{\partial\theta}\in\mathbb{R}^{3*3}$ (19)
When we consider a steady-state situation, i.e. $\dot{q},\ddot{q}=0$, to analyze the variation of the maximum force output by the robot at different positions, we assume that the three motors always output their maximum torque. Then, using Equation (18), we can calculate the maximum external force F that the robot can hold.
In Fig. S7c, different colors represent the variation in normalized output force in the Z-axis direction at different positions, when the torque output by the three motors is always the same. Here, red represents the maximum output force, while blue indicates the minimum output force. This visualization reflects how the robot’s maximum load-bearing capacity changes with the position of the top surface. Overall, the conclusion is that the load-bearing capacity is greater at higher planar heights compared to lower positions.
Method S2: Spherical Coordinate-based Workspace Evaluation
Besides the XYZ-Coordinates of the center $O_{R}$ of the top plane, the spatial pose of the top plane can also be expressed as [$\psi$, $\delta$, $OR_{Z}$], where $OR_{Z}$ refer to the height of the plane. When the input angles of the three legs fall within the $[30\degree,90\degree]$ range, we can plot the maximum tilt angles of the top plane at different heights pointing in different directions, as shown in Fig. S8. The blue dots in the Fig. S8 illustrate the possible tilt angles of the plane in various directions when the height of the plane center is within the range of different heights. The azimuthal angle in polar coordinates represents the tilt direction of the plane’s normal vector, while the radial coordinate signifies the tilt angle of the directional vector in that direction. We can observe that the tilt angle of the plane can reach up to $46\degree$ in specific directions and when the height of the center of the surface is $8cm$. When the height is between $8cm$to $10cm$, the maximum tilt angle of the plane in various directions appears to be relatively uniform.
Method S3: Multi-module Transfer Strategy
In Chapter 4.4, we discussed the strategy for object translation using a single module. This section extends the discussion to multi-module systems by detailing the strategies required for inter-module object translation. The basis of this approach lies in the coordinated lowering of Surface 1 while the object is on Surface 2. The transition of the object between the module surfaces is achieved by leveraging transfer strategy with Surface 2 until it securely lands on Surface 1. This process involves carefully balancing positioning, timing, and movement control to ensure stable translation of the object (see Algorithm S1). Key to this strategy is the real-time adjustment and monitoring of both modules to facilitate a smooth transition, setting the stage for subsequent transfers and maintaining the system’s readiness for ongoing operations.
Algorithm S1 Multi-module Translation: Translate object from Surface 2 to 1
1:procedure Multi-module Translation: ($X_{1ref},Y_{1ref}$)
2:     ##\#\ #Line 4-5: While on $surface_{2}$, $surface_{1}$ tilts and lowers as $surface_{2}$ translates, guiding the object onto $surface_{1}$.
3:     if $(x,y)\ in\ surface_{1}$ then
4:         $Surface_{1}\leftarrow\left[0,-\theta_{\beta},H_{init}-1\right]$
5:         $Surface_{2}\leftarrow translation\left(X_{2ref},Y_{2ref}\right)$triangleright $Algorithm_{1}$
6:     else
7:         ##\#\ #Line 8-9: Once the object fully moves to $surface_{1}$, execute the single-module transition algorithm, and reset $surface_{2}$.
8:         $Surface_{2}\leftarrow\left[0,0,H_{init}\right]$
9:         $Surface_{1}\leftarrow translation\left(X_{1ref},Y_{1ref}\right)$triangleright $Algorithm_{1}$
10:     end if
11:end procedure
Figure S7: Module design and workspace analysis. (a) Module design. Each module consists of an acrylic top surface, origami-based legs fabricated from FR4 and polyimide layers, BLDC motors for actuation, PLA connectors, and an acrylic base platform. (b) Robot structural schematic. (c) Cartesian analysis: $\theta\in[30\degree,90\degree]$. Dots highlight viable center positions. Different colors represent the variation in normalized output force in the Z-axis direction at different positions while maintaining uniform torque outputs across the three motors. Here, red represents the maximum output force, while blue indicates the minimum output force.
Figure S8: Workspace analysis in spherical coordinates. The maximum tiltable angle changes with elevation and direction. In the polar plots, azimuth signifies tilt direction, and radius quantifies potential tilt based on set conditions. The tilt corresponds to the plane’s normal vector in spherical coordinates.

### v2

@@ 4.3 Control System Overview and Implementation
The overall control of the developed system includes five main components (Fig. 2). Manipulation strategy transitions are managed by predefined task objectives. This system employs a state machine that dynamically alters between three modes: translation, rotation, and flipping. The position and the orientation of the object are obtained by the real-time sensor data from the vision system, guiding the dynamic generation of the surfaces’ roll, pitch, and height references in accordance with the appropriate manipulation strategy. Using the inverse kinematics algorithm explained in Supplementary Notes 1, the desired joint angles are computed. These reference values are then input into a PID-based servo motor controller, ensuring precise trajectory tracking of the joint commands. In the control scheme for the BLDC motor, as illustrated in Fig. 2c, a cascade control structure is implemented, comprising both a feedforward component and a current loop within the position loop.
@@
2:  #\#\ Line 3-4: compute the position error between the target and the object’s current position.
@@
6:   #\#\ Line 7-10: if the positional error exceeds $\epsilon$, calculate the reference pose for the surface.
7:   $\theta_{surface\_roll}=\textrm{PID}(Y_{error})$ $\triangleright$ Equation 2
8:   $\theta_{surface\_pitch}=\textrm{PID}(X_{error})$ $\triangleright$ Equation 2
9:   $H=A\sin(\omega t)+H_{init}$ $\triangleright$ Equation 3
@@
$\begin{split}Ref_{roll}&=K_{P}Y_{error}(t)+K_{i}\int_{0}^{t}Y_{error}(\tau)d\tau+K_{d}\frac{d}{dt}Y_{error},\\ Ref_{pitch}&=K_{P}X_{error}(t)+K_{i}\int_{0}^{t}X_{error}(\tau)d\tau+K_{d}\frac{d}{dt}X_{error}.\end{split}$ (2)
In low-friction scenarios, where the friction coefficient $\mu$ is smaller than $tan(\psi)$ (the tangent of the surface’s normal vector polar angle $\psi$), the control approach resembles a classic ball balancing table. Here, adjusting the surface’s inclination smoothly guides the object to its target. When the friction coefficient between two objects exceeds 0.33, inclination alone may not suffice. To overcome this, we modulate the surface’s height as well as its orientation (Fig. 2b), introducing vertical oscillations that reduce frictional resistance and apply periodic impulses to move the object. We define:
@@
where $t$ is the elapsed time and $\omega$ is the oscillation frequency, adjustable to suit different objects. The elevation of the surface also affects its maximum achievable tilt angle (see Supplementary Fig. 2). To maintain a large range of motion, here the vertical displacement amplitude spans from 8 cm to 10 cm. Friction affects manipulation by influencing the required oscillation amplitude and movement speed, creating a trade-off between operational speed and the safe handling of delicate objects.
@@
$C_{WR}=\begin{bmatrix}&\cos(\theta)\ &0\ &sin(\theta)\\ &\sin(\phi)\sin(\theta)\ &\cos(\phi)\ &-\cos(\theta)\sin(\phi)\\ &-\cos(\phi)\sin(\theta)\ &\sin(\phi)\ &\cos(\phi)\cos(\theta)\\ \end{bmatrix}$. (4)
@@
$N_{O_{R}}=C_{WR}*[0,0,1]^{T}=[\sin(\theta),-\cos(\theta)\sin(\phi),\cos(\phi)\cos(\theta)]^{T}.$ (5)
@@
$\begin{split}\delta&=arctan(\frac{N_{O_{R}}(y)}{N_{O_{R}}(x)})=\arctan(\frac{-\cos(\theta)\sin(\phi)}{\sin(\theta)}),\\ \psi&=\arccos(N_{O_{R}}(z))=\arccos(\cos(\phi)\cos(\theta)).\end{split}$ (6)
Once we determine the required orientation and height parameters $(\delta,\psi,H)$ for the Canfield mechanism’s end-effector, we apply inverse kinematics (Supplementary Notes 1) to compute the corresponding joint angles.
@@
2:  #\#\ Line 3-4: initialize the surface.
@@
5:  #\#\ Line 6-21: repeat when the angular error between the object’s current angle and target angle exceeds $\epsilon$. This constitutes one cycle.
@@
7:   $\theta_{surface2\_roll}=\textrm{PID}(\theta_{ref}-\theta_{yaw})$ $\triangleright$ Equation 7
@@
13:   #\#\ rotate the surface to the reference pose to induce the object’s rotation.
@@
16:   delay 500msdelay\ 500ms
17:   #\#\ reset the surface after one cycle to establish new contact points in next cycle.
@@
22:  #\#\ Line 23-24: reset the surface once the object reaches the reference angle to prepare for the next task.
@@
$\theta_{surface2\_roll}=K_{P}*\theta_{error}(t)+K_{i}\int_{0}^{t}\theta_{error}(\tau)d\tau+K_{d}\frac{d}{dt}\theta_{error}.$ (7)
@@
2:  #\#\ Line 3-5: based on the object’s size, transport it from the initial position to a flipping position between the two surfaces.
@@
4:   $Surface_{1}\leftarrow\textrm{translation}\left(X_{ref},Y_{ref}\right)$ $\triangleright$ Algorithm 1
@@
7:   #\#\ Line 8-9: Lower the surface to initialize the flipping task.
@@
12:   #\#\ Quickly raise Surface 2 to generate an impact, flipping the object.
13:   $Surface_{2}\leftarrow\left[0,0,H_{high}\right]\ in\ 100ms$
14:   delay 1000msdelay\ 1000ms
15:   #\#\ Line 16-17: after flipping, reset the surface to prepare for the next task.
@@
$\begin{split}0-\frac{1}{2}J\omega_{t1}^{2}&\geq\frac{L}{2}mg(\sin(\theta)-1),\\ \omega_{t1\_min}&=\sqrt{\frac{mgL(1-\sin(\theta))}{J}}.\end{split}$ (8)
@@
Data availability
All data needed to evaluate the conclusions in the paper are present in the paper or the Supplementary Materials.
@@ Acknowledgements
The authors thank Alexander Schüßler of Reconfigurable Robotics Lab, EPFL for his contributions to using computer vision to identify the position of deformable objects. This research has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement no: 101069536 (MOZART project).
@@
Competing Interests
The authors declare no competing interests.